Find the maximum and minimum value of the expression x2+14x+9/x2+2x+3 

Grade:11

1 Answers

Vikas TU
7 years ago
Let N = x2+14x+9
and 
D = x2+2x+3 
Differntiating both we get,
dN/dx = 2x + 14 = 0 = > x = -7
and
dD/dx = 2x + 2 = 0 => x = -1
At x = -1 x2+14x+9/x2+2x+3  = > -2 
At x = – 7 x2+14x+9/x2+2x+3  = > 20/19.
Hence, -2 is the minimum value and 20/19 is the maximum value.